An object is thrown up at angle of θ=30º with the initial speed of 19m/s.

i) Find the maximum height of the object and the maximum horizontal distance.

ii) Find the minimum speed of the object during the motion.

iii) Find the total time in the air.
Answer #1

A)

Max height, H = (u sin(theta))^2/2g

H = (19 sin(30))^2/2 x 9.8 = 4.60 m

Range, R = u^2 sin(2 theta)/g

R = 19^2 sin(60)/9.8 = 31.90 m

B)

Minimum Velocity is at max height

V = 19 cos 30 = 16.454 m/s

C)

Time, t = 2u sin(theta)/g

t = 2 x 19 sin(30)/9.8 = 1.94 sec

Comment in case any doubt please rate my answer.....